# The Evolution of Reinforcement Learning in Quantitative Finance

### v1

@@ Abstract.
copyright: acmcopyrightjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Applied computing Electronic commerceccs: Information systems Expert systems
@@ 2.2. Sample Efficiency
A hybrid approach can be beneficial. For example, an RL agent could be pre-trained using off-policy methods on historical data to develop a baseline strategy. This agent can then be fine-tuned using on-policy methods in a live environment for stability and adaptability. In practice, such as in equity trading or broader quantitative finance setups, an agent initially trained with DQN can switch to PPO in live trading to balance stability and responsiveness, combining the benefits of both methods (RAFAILOV_2023, ; ZHAO_2023, ). General care should be taken when using historical data, as it could lead to biases and unintended trading behaviours. In portfolio management, reliance on past data can result in strategies that perform poorly under new market conditions due to non-stationary dynamics (LIANG_2018, ). Thorough testing and validation are necessary to mitigate these risks.
@@ 3. Main Reinforcement Learning Methods
Table 1 presents a classification of each publication reviewed in this survey according to the RL method and algorithm used. Within the Value-based approach, DQN and Q-learning (WATKINS_1989, ) emerge as the most frequently used algorithms. Recurrent Reinforcement Learning (RRL) (MOODY_1997, ; MOODY_1998, ; MOODY_1998_2, ; MOODY_2001, ) assumes dominance in the Policy-based methodology, and PPO (SCHULMAN_2017, ), and DDPG (LILLICRAP_2015, ) are the most popular in the Actor-Critic approach. Overall, the Value-based method has received the most extensive research, followed by the Policy-based, Actor-Critic, and model-based methods. In the interest of simplicity, we have occasionally categorised publications under the “other” bucket when the proposed algorithm is non-standard or infrequently used. Furthermore, we have classified several publications under multiple categories due to their use of a diverse range of algorithms. Lastly, we have refrained from including publications that fall outside these four categories for the sake of brevity. For example, (YANG_2018, ) applies Inverse RL (IRL) (NG_2000, ) to QF, a method that is not discussed in this survey.
@@ 3.2. Policy-Based Method
Policy-based methods represent the second most explored area of RL in the reviewed literature. These methods provide a direct mapping from states to actions, eliminating the need to compute the expected outcome of different actions as in the Value-based approach, resulting in faster learning processes.
A significant advantage of Policy-based methods is the continuous action space for the agent. Consider a portfolio of stocks: with the Value-based approach, portfolio weights can only take discrete values like buy, sell, or hold. In contrast, the Policy-based approach allows portfolio weights to assume any value in $\left[0,1\right]$ in the long-only case.
@@ 3.4. Model-Based RL
Model-based methods in RL offer several advantages and present unique challenges, particularly when applied to financial trading systems 666see also Section 2.:
@@ 4.1. Introduction
Financial markets epitomise a PO environment, lacking deterministic representation; a single day’s pricing cannot encapsulate the state of the environment, unlike full observability in Atari video games (MNIH_2013, ; MNIH_2015, ). The use of LSTM and RNN models is particularly relevant in this context, as they are capable of capturing long-term dependencies and patterns in sequential data. These models help address the limitations of the Markov assumption by incorporating memory effects and providing a more robust representation of the financial market dynamics.
In early adoptions of RL in QF, the environment was assumed to be fully observable (NEUNEIER_1996, ), mainly because the inherent randomness of financial markets makes an exact environment representation virtually impossible. Furthermore, adding more features does not necessarily improve performance (MARTINEZ_2009, ), necessitating strategic feature selection to manage randomness, avoid the curse of dimensionality (BELLMAN_1957, ), and address interpretability issues. Finally, we note that empirical studies have shown that models that incorporate memory effects, such as LSTM and RNN, significantly outperform those based on the Markov assumption, particularly in predicting long-term trends and capturing market anomalies (FISCHER_2018_2, ).
@@ 4.2. Features
Modelling financial markets with their inherent randomness is a considerable challenge. Consequently, input data, feature selection, and extraction become critical to the success of an RL system. The literature presents a diverse range of data sources, features, and feature selection mechanisms, reflecting the current state of research and highlighting the domain-specific relevance of different features. Table 2 presents features and data used in the top fifteen most cited papers in our sample (as of October 2022), which are also ordered chronologically.
Publication Data Freq. Period PH TI CP ME PV CH OB Oth. FM
Neuneier (NEUNEIER_1996, ) Daily 1986-1996 x x x x
Moody and Wu (MOODY_1997, ) Monthly/30-min. Several x x
Moody et al. (MOODY_1998, ) Monthly/30-min. Several x x
Moody and Safell (MOODY_2001, ) Monthly/30-min. Several x x
Dempster et al. (DEMPSTER_2001, ) 1-min. 1994-2002 x x x
Dempster and Leemans (DEMPSTER_2006, ) 1-min. 2000-2002 x x x
Lee et al. (LEE_2007, ) Daily 1999-2005 x x x x
Deng et al. (DENG_2016, ) Tick-level Several x x x
Jiang and Liang (JIANG_2017, ) 30-min. 2014-2017 x x x x
Jiang and Liang (JIANG_2017_1, ) 30-min. 2014-2017 x x x x
Almahdi and Yang (ALMAHDI_2017, ) Weekly 2011-2015 x x
Liang et al. (LIANG_2018, ) Daily 2015-2017 x x x
Xiong et al. (XIONG_2018, ) Daily 2009-2018 x x x x
Jeong and Kim (JEONG_2019, ) Daily Several x x
@@
Technical analysis consists of indicators or rules designed to predict the direction of the price using past prices and volume variables. However, the effectiveness of these indicators is often challenged by the Efficient Market Hypothesis (EMH), which posits the unpredictability of stock market prices (FAMA_1970, ). As such, research on these indicators has produced mixed results, and some studies have methodological issues in their testing procedures (PARK_2007, ).
Despite such criticism, numerous academics have questioned the EMH, arguing that future prices can be predicted based on past prices, thus challenging the random walk hypothesis (LO_1988, ). Consequently, several researchers have used technical indicators to represent the agent’s environment within this domain, including (DEMPSTER_2001, ; DEMPSTER_2002, ; DEMPSTER_2006, ; PARK_2020, ; WU_2020, ; ZHANG_2020_A, ; WANG_2021_B, ). However, there are also publications that question the efficacy of technical analysis in the realm of RL in QF (DEMPSTER_2006, ).
Key technical indicators often used include the Moving Average (MA), Exponential Moving Average (EMA), Moving Average Convergence/Divergence (MACD), Japanese candlestick, and the Relative Strength Index (RSI) (MURPHY_1999, ). Moreover, some researchers have created features based on the original definitions of technical indicators, demonstrating the adaptability of these tools (GABRIELSSON_2015, ).
@@
Accounting data from financial statements underpin traditional factor investing, with numerous strategies developed over half a century based on such data aimed at explaining expected asset returns (COCHRANE_2005, ). As identified by Harvey and Liu (HARVEY_2019, ), more than 400 factors have been proposed, covering a wide spectrum from macroeconomic to statistical commonalities. With advances in computational power, the discovery of these factors and their risk premiums has been streamlined. Various research studies, for example, (FAMA_1992, ), have revealed factors that explain equity returns and contribute to long-term active outperformance. Prominent factors include Price momentum [^8] Trading strategies where one buys winner stocks that performed well and sells those that did poorly in the past to generate profits over a 3-12 months holding horizon. (JEGADEESH_1993, ), Value [^9] Buying stocks with relatively low prices based on fundamental ratios and following the opposite when the high share prices are detected. (FAMA_1992, ), Size [^10] Small companies perform better than big companies based on some measure, such as market capitalisation. (BANZ_1981, ), Quality [^11] Buying good quality companies and selling poor quality companies based on information reported in the financial statements. (GRAHAM_1973, ), Low beta [^12] Buying stocks with low risk and selling stocks with high risk. (FRAZZINI_2014, ). These factors can be employed as features to describe the RL agent’s environment.
These features are not often used in the RL-based financial literature, likely due to the infrequent updates of accounting data, which occur quarterly. This temporal mismatch, as in the case of the price-to-earnings ratio, can limit the effectiveness of these factors as daily signals. However, some researchers have begun to integrate accounting data or investment factors derived from RL. Zhang and Maringer (ZHANG_2013, ; ZHANG_2016, ) pioneered the use of accounting ratios in an evolutionary feature selection scheme, while Wang et al. (WANG_2019, ) included ratios like price to earnings in state representations. More recently, Coqueret and André (CORQUERET_2022, ) demonstrated the effective incorporation of various investment factors, including those rooted in accounting data, into RL-based trading systems.
@@
Among the literature surveyed, we find a substantial focus on the areas of trade execution [^13] The agent (for example, a broker) receives and completes an order either as a buy or sell with a given quantity. and market making in financial markets. These specific applications tend to employ a unique set of features to model the environment, distinguishing them from other domains. While we explore these applications more comprehensively in subsections 7.4 and 7.5, it is beneficial to highlight some of the most commonly observed features in the associated literature. For example, Chan and Shelton (CHAN_2001, ) used order imbalance-based features [^14] Order imbalance is when an excess of buy or sell orders for an asset causes difficulties in matching the orders of buyers and sellers., depth of the market [^15] The amount of price change until the order is filled., and time to fill a limit order, among others. Other authors used the Bid-Ask spread [^16] The amount by which the Ask price exceeds an asset’s Bid price., and remaining inventory, among others, as part of their environment (NING_2020, ). Numerous other features are also deployed in this context, but a detailed discussion is omitted here for brevity.
@@
The impact of sentiment data on stock prices is well documented in the financial literature (ANTWEILER_2004, ; TETLOCK_2007, ). Its application in RL strategies has also been demonstrated (FEUERRIEGEL_2016, ; KAUR_2017, ; YANG_2018, ; NAN_2020, ; YE_2020, ; SAWHNEY_2021, ; KORATAMADDI_2021, ; HUANG_2022, ). Feuerriegel and Prendinger (FEUERRIEGEL_2016, ) pioneered the use of sentiment data, employing DGAP for an RL-based trading strategy. Kaur (KAUR_2017, ) followed, enriching the environment of the Q-learning agent with sentiment data. In particular, diverse sources such as Reuters News Corpus, Twitter, Google News, and Thomson Reuters News analytics have been used to create sentiment signals for RL strategies, affirming that diverse information mitigates uncertainty in the environment (YANG_2018, ; NAN_2020, ; YE_2020, ; SAWHNEY_2021, ; KORATAMADDI_2021, ). However, the use of sentiment/news data presents challenges. It primarily covers prominent companies, limiting its utility to lesser-known stocks. Despite sentiment’s potential value, its integration into state representation remains scarce in published work.
@@
Macroeconomic data plays a pivotal role as it encapsulates the overarching economic landscape in which companies operate. This encompasses widely recognised indicators such as interest rates, inflation, and GDP. In particular, cyclical companies are intimately tied to these macroeconomic variables due to their strong correlation with the state of the economy. Despite the importance of macroeconomic data, its incorporation into the state representation is found in a limited number of studies in the existing literature. Neuneier’s work (NEUNEIER_1996, ) is a likely pioneer in this regard, incorporating interest rates into the analysis (FISCHER_2018, ). Subsequently, Benhamou’s series of studies (BENHAMOU_2020_A, ; BENHAMOU_2020_B, ; BENHAMOU_2021_A, ) further integrated the slope of the US treasury yield curve, serving as a precursor risk indicator within contextual information.
@@
Earlier subsections discussed numerous environmental features implemented or potentially applicable within finance-related RL scenarios. The selection and use of these features often depends on the specific application, resulting in bespoke adaptations. This section consolidates such unique variations.
In trade execution applications, state representation often incorporates aspects such as the elapsed time, the remaining shares to order (NEVMYVAKA_2006, ), and the current time (NING_2020, ). Deng et al. (DENG_2015, ) employ over 80 raw features, including non-standard ones like the volume-weighted average price (VWAP), historical bond equity correlations and credit spreads of global corporate bonds (BENHAMOU_2020_A, ; BENHAMOU_2020_B, ; BENHAMOU_2021_A, ; BENHAMOU_2021_B, ). Other studies integrate current and cash positions into state representation (SHERSTOV_2004, ; BERTOLUZZO_2012, ; JIN_2016, ; JIANG_2017, ; JIANG_2017_1, ; KAUR_2017, ). The financial literature and the interest of practitioners have spotlighted several underexplored datasets in RL in finance, including the following.
  1. (1)
    Environmental, Social, and Governance (ESG) factors have emerged as significant themes in the investment world (GOMPERS_2003, ; OIKONOMOU_2012, ) recently. The integration of ESG-based features into the RL environment may pave the way for novel research trajectories.
  2. (2)
    Supply chain data, a valuable dataset in finance literature (COHEN_2008, ; MENZLY_2010, ), offers insights into the interconnected relationships among firms, including suppliers, customers, competitors, and joint ventures. These relationships, varying in transparency, result in unstructured, incomplete, and highly complex supply chain data. Integrating supply chain data into the RL environment could help capture the intricate market dynamics of individual stocks in relation to others.
@@ 4.3. Feature Selection and Feature Extraction
As outlined in subsection 4.2, the complexity of financial markets requires many features to define an agent’s environment. However, this proliferation of features inflates the dimensionality of the environment and requires intense computational resources. During the early years of Value-based applications when computation was more of a limiting factor, authors extensively discretised the state to mitigate this issue, for instance, Dempster and Romahi’s work (DEMPSTER_2002, ) where they binarised eight technical indicators, forming a bit string (HOLLAND_1992, ). A significant change in the feature selection and representation framework is observed after 2016, mainly due to advances in DRL (MNIH_2013, ; MNIH_2015, ) and the adoption of NN architectures, such as Convolutional Neural Networks (CNN) (LECUN_1995, ; KRIZHEVSKY_2012, ), propelled by improved computational capabilities. This subsection discusses the evolving mechanisms for feature selection and extraction in financial RL applications.
4.3.2. Genetic Algorithm as a Tool for Feature Selection.
The extensive use of Genetic Algorithms (GA) in RL applications traces back to Dempster and Romahi (DEMPSTER_2002, ), who proposed a hybrid evolutionary RL system with GA used for optimal feature selection. Bates et al. (BATES_2003, ) extended this framework to include order flow and order book features in addition to technical indicators. Hryshko and Downs (HRYSHKO_2004, ) mirrored this approach but suggested the Stirling ratio as the fitness function, reducing computational demands by removing the role of the RL algorithm in determining the count of state representation features.
Chen et al. (CHEN_2007, ) proposed a Genetic Network Programming method based on SARSA to select relevant technical indicators from a candlestick chart  [^17] A candlestick chart visually represents an asset’s price movements over a specific period, showing open, close, high, and low prices., a strategy also adopted by Mabu et al. (MABU_2007, ) within an Actor-Critic framework. Gu et al. (GU_2011, ) further developed Chen et al.’s work by incorporating plural subroutines. Zhang and Maringer (ZHANG_2013, ) introduced GA as a pre-screening tool, feeding selected features into an RRL trading system. The model encompassed eight Boolean indicator variables, sourced primarily from technical analysis. Selection was based on an average fitness function scheme, providing an improvement over Maringer and Ramtohul (MARINGER_2012, ) by reducing overfitting. Zhang and Maringer (ZHANG_2016, ) later refined their approach employing a roulette wheel selection approach during the training phase. Zhu et al. (ZHU_2018, ) applied GA to choose features for a Support Vector Regression model that predicts stock trading bounds based on Q-learning.
@@
Earlier, we mentioned that the introduction of DL and DRL had significantly impacted the RL literature. From 2016 onwards, many publications in this survey have used deep network architectures. Therefore, in this section, we will highlight some key advancements instead of listing all the developments in the literature. A critical paper on current RL-based trading systems is by Deng et al. (DENG_2016, ). The authors presented the first comprehensive framework, namely DRRL, that uses DL as a feature extraction tool and applies the RRL framework. Moreover, the authors propose “task-aware backpropagation thought time” (task-aware BPTT) to solve the gradient vanishing problem (BENGIO_1994, ).
Recurrent Neural Network and Long Short-Term Memory: The RRL and its derivative DRRL are based on Recurrent Neural Networks (RNN), a vital tool in POMDPs due to their ability to record and use past states to optimise future actions (RUMELHART_1985, ; DENG_2016, ; GERS_2000, ). However, vanishing gradients in deep structures is an issue (BENGIO_1994, ). Hence, Lu (LU_2017, ) adopted Long Short-Term Memory (LSTM) networks (HOCHREITER_1997, ) with dropout incorporated (GOODFELLOW_2016, ) for feature learning and capturing the underlying financial market conditions.
@@
The use of LSTM in decision-making is advantageous due to its ability to remember features over longer steps, an essential quality for financial market trading. In addition, its memory cell retains trading actions, thereby capturing the impact of transaction costs more effectively. Bisht and Kumar (BISHT_2020, ) developed two LSTM-based systems, one functioning as an autoencoder for feature extraction and the other using the decoded features within an LSTM module for decision-making. Lin and Beling contrasted two network structures, one combining a fully connected network (FCN) and LSTM and the other based purely on FCN (LIN_2020, ). They concluded that the former generally outperformed the benchmarks. However, regular LSTM’s assumption that information is uniformly spaced can be problematic in varying data environments. To address this, Sawhney et al. proposed a time-aware LSTM (SAWHNEY_2021, ; BAYTAS_2017, ). Other notable applications of LSTM can be found, for example, in (CHEN_2019, ; JIA_2019, ; PONOMAREV_2019, ; TSANTEKIDIS_2020, ; KUMAR_2022, ).
Gate Recurrent Units (GRU) (CHO_2014, ) are used to distill informative characteristics from raw financial data (WU_2020, ; LIU_2020, ), as studies such as Chung et al. (CHUNG_2014, ) have shown their superiority over LSTM in handling time series data.
The Advent of Convolutional Neural Networks: In the current RL financial literature, CNNs are commonly employed for feature extraction. Jiang et al. (JIANG_2017, ; JIANG_2017_1, ) were among the first to apply CNN and other network architectures, specifically in the context of PM. They introduced an Ensemble of Identical Independent Evaluators, where the price movement of each asset is extracted independently. However, as the network depth increases, this CNN structure could face challenges such as vanishing gradient and gradient explosion. Additionally, this approach may fail to capture any non-linear relationships among assets. Consequently, Liang et al. (LIANG_2018, ) adopted a similar framework but replaced CNN with a Deep Residual Network (HE_2016, ).
Shi et al. (SHI_2019, ) highlight a gap in RL-based PM literature, which primarily relies on price-based features, neglecting temporal information across different scales. To address this, they proposed the Ensemble of Identical Independent Inception using the Inception Network for simultaneous multiscale application (SZEDEGY_2015, ). Their innovative architecture integrates three parallel components: short-term and medium-term price movement features and maximum price features. This pioneering work demonstrates the value of parallel, multi-component feature extraction in network architectures.
Soleymani and Paquet (SOLEYMANI_2021, ) apply an autoencoder for feature extraction and dimensionality reduction. A Graph Convolutional Network (HENAFF_2015, ) is then used to capture the interactions among financial assets. Aboussalah et al. (ABOUSSALAH_2022, ) implemented dual CNNs in an Actor-Critic agent, with one for time series and the other for cross-sectional features, the first attempt to combine these features. We note that this is the first study to combine time series and cross-sectional features. However, the results may exhibit survivorship bias due to stock selection. Some other publications using CNNs include (TAGHIAN_2022, ; BENHAMOU_2020_A, ; BENHAMOU_2020_B, ; BENHAMOU_2021_A, ; BENHAMOU_2021_B, ).
Introducing Attention Mechanisms: Several network architectures have been used as feature mechanisms across a range of applications. However, effectively estimating the temporal relationships between assets remains a challenge. To address this, the integration of an attention layer (SUTSKEVER_2014, ; VASWANI_2017, ) in the network architecture has been proposed to capture temporal and time series data. Consequently, variations in attention have garnered considerable interest within QF in the RL framework.
Lei et al. (LEI_2020, ) employed a GRU to model long-term relationships, complemented by a temporal attention mechanism (PEI_2017, ) to discern stock interdependencies. Weng et al. (WENG_2020, ) proposed the use of separable convolution (CHOLLET_2017, ) and three-dimensional gated attention networks for feature extraction. Zhang et al. (ZHANG_2020_B, ) took a unique approach to model asset interdependencies, introducing a two-module network; the first was an LSTM network capturing sequential asset patterns, while the second employed dilated causal convolutions (OORD_2016, ) and correlational convolutions for asset correlations. The output of both modules is concatenated for comprehensive analysis. Further studies using attention mechanisms are presented in (WANG_2019, ; WANG_2021_A, ; XU_2021, ).
Other Deep Learning-Based Feature Mechanisms: As discussed in Section 1, a significant challenge for RL is the need for extensive financial datasets. Addressing this, Yu et al. (YU_2019, ) proposed a data augmentation mechanism within their model-based Actor-Critic framework, using a Generative Adversarial Network (GAN) (GOODFELLOW_2014, ). Specifically, they adopted a Recurrent GAN (RGAN) (ESTEBAN_2017, ) variant, where the generator and discriminator are replaced by RNNs. They employed an RGAN to produce asset closing price returns at a frequency higher than the RL agent, which was then downsampled for synthetic High-Low-Close data. Moreover, AbdelKawy et al. (ABDELKAWY_2021, ) used a Deep Belief Network (DBN) (HINTON_2006, ) for feature extraction and dimensionality reduction.

### v2

@@ Abstract.
copyright: acmcopyrightjournalyear: 2025doi: XXXXXXX.XXXXXXXjournal: CSURjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Computing methodologies Machine learningccs: Applied computing Electronic commerceccs: Information systems Expert systems
@@ 3. Main Reinforcement Learning Methods
Table 1 presents a classification of each publication reviewed in this survey according to the RL method and algorithm used. Within the Value-based approach, DQN and Q-learning (WATKINS_1989, ) emerge as the most frequently used algorithms. Recurrent Reinforcement Learning (RRL) (MOODY_1997, ; MOODY_1998, ; MOODY_1998_2, ; MOODY_2001, ) assumes dominance in the Policy-based methodology, and PPO (SCHULMAN_2017, ), and DDPG (LILLICRAP_2015, ) are the most popular in the Actor-Critic approach. Overall, the Value-based method has received the most extensive research, followed by the Policy-based, Actor-Critic, and model-based methods. In the interest of simplicity, we have occasionally categorised publications under the “other” bucket when the proposed algorithm is non-standard or infrequently used. Furthermore, we have classified several publications under multiple categories due to their use of a diverse range of algorithms and refrained from including publications that fall outside these four categories.
@@ 3.2. Policy-Based Method
Policy-based methods represent the second most explored area of RL in the reviewed literature. These methods provide a direct mapping from states to actions, eliminating the need to compute the expected outcome of different actions as in the Value-based approach, resulting in faster learning processes. A significant advantage of Policy-based methods is the continuous action space for the agent. Consider a portfolio of stocks: with the Value-based approach, portfolio weights can only take discrete values like buy, sell, or hold. In contrast, the Policy-based approach allows portfolio weights to assume any value in $\left[0,1\right]$ in the long-only case.
@@ 3.4. Model-Based RL
Model-based methods in RL offer several advantages and present unique challenges, particularly when applied to financial trading systems:
@@ 4.1. Introduction
Unlike full observability in video games (MNIH_2013, ; MNIH_2015, ), financial markets epitomise a PO environment, lacking deterministic representation; a single day’s pricing cannot encapsulate the state of the environment. The use of LSTM and RNN models is relevant here, as they can capture long-term dependencies and patterns in sequential data. These models help address the limitations of the Markov assumption by incorporating memory effects and providing a dynamic representation of the financial markets.
In early adoptions of RL in QF, the environment was assumed to be fully observable (NEUNEIER_1996, ), mainly because the inherent randomness of financial markets makes an exact environment representation impossible. Adding more features does not necessarily improve performance (MARTINEZ_2009, ), necessitating strategic feature selection to manage randomness, avoid the curse of dimensionality (BELLMAN_1957, ), and address interpretability issues. We note that empirical studies have shown that models that incorporate memory effects, such as Long Short-Term Memory (LSTM) (HOCHREITER_1997, ) and Recurrent Neural Networks (RNN) (RUMELHART_1985, ; DENG_2016, ; GERS_2000, ), significantly outperform those based on the Markov assumption, particularly in predicting long-term trends and capturing market anomalies (FISCHER_2018_2, ).
@@ 4.2. Features
Financial market modelling poses significant challenges due to inherent randomness. Thus, the choice and extraction of input data and features are crucial for the effectiveness of RL systems. The literature showcases a variety of data sources and feature selection methods that underline their specific relevance within this domain. Table 2 (in Appendix) details the features and data from the top fifteen most-cited papers in our sample as of October 2022, listed chronologically.
@@
Technical analysis employs indicators and rules to predict price directions based on past price and volume data (MURPHY_1999, ). The validity of these methods is frequently contested by the Efficient Market Hypothesis (EMH), which posits that stock prices are inherently unpredictable (FAMA_1970, ). Despite mixed research outcomes and some methodological criticisms (PARK_2007, ), numerous scholars challenge the EMH by suggesting that past prices can forecast future trends, thereby refuting the random walk hypothesis (LO_1988, ). Consequently, technical indicators such as Moving Average (MA), Exponential Moving Average (EMA), Moving Average Convergence/Divergence (MACD), Japanese candlestick, and Relative Strength Index (RSI) are widely used to represent environments within Reinforcement Learning (RL), as evidenced in various studies (e.g., (DEMPSTER_2001, ; DEMPSTER_2002, ; DEMPSTER_2006, ; GABRIELSSON_2015, ; PARK_2020, ; WU_2020, ; ZHANG_2020_A, ; WANG_2021_B, )). However, their efficacy in RL applications remains a subject of debate (DEMPSTER_2006, ).
@@
Accounting data underpins traditional factor investing, with numerous strategies (HARVEY_2019, ) developed over decades to explain expected asset returns (COCHRANE_2005, ). Key factors include price momentum [^7] Trading strategies where one buys winner stocks that performed well and sells those that did poorly in the past to generate profits over a 3-12 months holding horizon. (JEGADEESH_1993, ), value [^8] Buying stocks with relatively low prices based on fundamental ratios and following the opposite when the high share prices are detected. (FAMA_1992, ), size [^9] Small companies perform better than big companies based on some measure, such as market capitalisation. (BANZ_1981, ), quality [^10] Buying good quality companies and selling poor quality companies based on information reported in the financial statements. (GRAHAM_1973, ), and low beta [^11] Buying stocks with low risk and selling stocks with high risk. (FRAZZINI_2014, ). Despite their proven efficacy, these factors are infrequently used in RL due to the quarterly updates of accounting data, limiting their effectiveness as daily signals. Nevertheless, pioneering work by Zhang and Maringer (ZHANG_2013, ; ZHANG_2016, ) and recent efforts by Wang et al. (WANG_2019, ) and Coqueret and André (CORQUERET_2022, ) have begun integrating these factors into RL environments, potentially enhancing the models’ predictive capabilities.
@@
Among the literature surveyed, we find a substantial focus on the areas of trade execution and market making. These specific applications tend to employ a unique set of features to model the environment, distinguishing them from other domains. While we explore these applications more comprehensively in subsections 7.4 and 7.5, it is beneficial to highlight some of the most commonly observed features in the associated literature. For example, Chan and Shelton (CHAN_2001, ) used order imbalance-based features, depth of the market, and time to fill a limit order, among others. Other authors used the Bid-Ask spread and remaining inventory, among others, as part of their environment (NING_2020, ). Numerous other features are also deployed in this context, but a detailed discussion is omitted here for brevity.
@@
The impact of sentiment data on stock prices is well documented in the financial literature (ANTWEILER_2004, ; TETLOCK_2007, ). Its application in RL strategies has also been demonstrated (FEUERRIEGEL_2016, ; KAUR_2017, ; YANG_2018, ; NAN_2020, ; YE_2020, ; SAWHNEY_2021, ; KORATAMADDI_2021, ; HUANG_2022, ). Diverse sources like Reuters News Corpus, Twitter, and Thomson Reuters News analytics have proven effective in creating sentiment signals that reduce uncertainty in financial models. However, the application of sentiment data in RL is predominantly focused on well-known companies, thus limiting its broader applicability to lesser-known stocks. Despite its demonstrated value, sentiment data is underutilised in state representations across the existing RL-based financial models.
@@
Macroeconomic data, including widely recognised indicators such as interest rates, inflation, and GDP, plays a crucial role by reflecting the broader economic context in which companies operate, particularly for cyclical firms closely aligned with economic fluctuations. However, the integration of such data into state representations in RL models is rare. Neuneier’s work (NEUNEIER_1996, ) was pioneering in incorporating interest rates (FISCHER_2018, ), and Benhamou’s subsequent studies (BENHAMOU_2020_A, ; BENHAMOU_2020_B, ; BENHAMOU_2021_A, ) expanded this approach by including the slope of the US Treasury yield curve as a risk indicator.
@@
In earlier subsections, we covered several financial features which can be or are used in the current RL literature. The selection and use of these features often depend on the specific application, resulting in bespoke adaptations. This section consolidates a few of those unique variations. In trade execution, typical state representations include elapsed time, remaining shares to order (NEVMYVAKA_2006, ), and current time (NING_2020, ). Deng et al. (DENG_2015, ) used over 80 raw features, such as the volume-weighted average price (VWAP) and credit spreads of global corporate bonds (BENHAMOU_2020_A, ; BENHAMOU_2020_B, ; BENHAMOU_2021_A, ; BENHAMOU_2021_B, ). Other studies have incorporated current and cash positions into their models (SHERSTOV_2004, ; BERTOLUZZO_2012, ; JIN_2016, ; JIANG_2017, ; JIANG_2017_1, ; KAUR_2017, ). Additionally, emerging data sources such as Environmental, Social, and Governance (ESG) factors (GOMPERS_2003, ; OIKONOMOU_2012, ) and supply chain data (COHEN_2008, ; MENZLY_2010, ) could be potential candidates for integration into RL environments to enrich the understanding of market dynamics and firm interconnections.
@@ 4.3. Feature Selection and Feature Extraction
The complexity of financial markets necessitates numerous features to define an agent’s environment, which can increase dimensionality and computational demands. Early Value-based applications addressed (DEMPSTER_2002, ) this by discretising states, such as binarising technical indicators into bit strings (HOLLAND_1992, ). Significant advancements in feature selection and representation occurred after 2016, driven by improvements in DRL and the adoption of neural network (NN) architectures such as Convolutional Neural Networks (CNN) (MNIH_2013, ; MNIH_2015, ; LECUN_1995, ; KRIZHEVSKY_2012, ). We discuss here the mechanisms for feature selection and extraction in financial RL.
@@
RNNs are a vital tool in POMDPs due to their ability to record and use past states to optimise future actions (RUMELHART_1985, ; DENG_2016, ; GERS_2000, ). However, vanishing gradients in deep structures are an issue (BENGIO_1994, ). Consequently, many researchers have shifted focus to LSTMs in the reviewed literature. The use of LSTM in decision-making is advantageous due to its ability to remember features over longer steps, an essential quality for financial market trading. Additionally, its memory cell retains trading actions, thereby capturing the impact of transaction costs more effectively. For example, Bisht and Kumar (BISHT_2020, ) developed two LSTM-based systems, one functioning as an autoencoder for feature extraction and the other using the decoded features within an LSTM module for decision-making. Other notable applications of LSTM include those found in (LU_2017, ; CHEN_2019, ; JIA_2019, ; PONOMAREV_2019, ; TSANTEKIDIS_2020, ; LIN_2020, ; SAWHNEY_2021, ; KUMAR_2022, ).
Transitioning from RNNs and LSTMs, CNNs introduce a different paradigm in NN architecture. CNNs are a crucial aspect of feature extraction in PM due to their ability to independently analyse asset price movements and manage complex data structures. Early implementations (JIANG_2017, ; JIANG_2017_1, ) introduced the Ensemble of Identical Independent Evaluators, which independently extracted the price movement of each asset. Challenges such as vanishing gradients and the inability to capture non-linear relationships prompted a transition towards Deep Residual Networks (HE_2016, ) for enhanced robustness (LIANG_2018, ). Further advancements addressed the lack of temporal dynamics in financial models by employing the Inception Network for multi-scale analysis (SZEDEGY_2015, ), integrating multiple data aspects into a cohesive framework (SHI_2019, ). Additionally, innovations like the integration of autoencoders and dual CNNs have expanded boundaries by enhancing dimensionality reduction and combining time series with cross-sectional data analysis in an Actor-Critic framework, marking significant progress in the domain (SOLEYMANI_2021, ; ABOUSSALAH_2022, ). These applications underscore the evolving capabilities of CNNs to provide comprehensive insights into financial asset interactions. Further contributions to the CNN literature are evidenced by additional studies (TAGHIAN_2022, ; BENHAMOU_2020_A, ; BENHAMOU_2020_B, ; BENHAMOU_2021_A, ; BENHAMOU_2021_B, ).
The integration of attention mechanisms marks yet another advance in NN technology, significantly enhancing NN in QF by improving temporal relationship analysis and long-term dependency modelling (SUTSKEVER_2014, ; VASWANI_2017, ; PEI_2017, ). These mechanisms have been successfully applied in various settings to boost the performance of financial models, as evidenced by their use in diverse applications across recent studies (LEI_2020, ; WENG_2020, ; ZHANG_2020_B, ; WANG_2019, ; WANG_2021_A, ; XU_2021, )
Lastly, innovative approaches like Generative Adversarial Networks (GANs) (GOODFELLOW_2014, ; ESTEBAN_2017, ), Deep Belief Networks (DBNs) (HINTON_2006, ), and Gate Recurrent Units (GRU) (CHO_2014, ) have been used in the reviewed literature for data augmentation (YU_2019, ) and for feature extraction are addressing the scarcity of extensive financial datasets (ABDELKAWY_2021, ; WU_2020, ; LIU_2020, ).
@@
Genetic Algorithms (GA) have been widely used in RL for different tasks. These include selecting the best features, optimising how models are evaluated, using extra data like order flow and order books, choosing the right technical indicators, and improving trading strategies. Several studies have explored these uses (e.g., (DEMPSTER_2002, ; BATES_2003, ; HRYSHKO_2004, ; CHEN_2007, ; MABU_2007, ; GU_2011, ; ZHANG_2013, ; MARINGER_2012, ; ZHANG_2016, ; ZHU_2018, )).